OpenStudy (anonymous):

The graph of this quadratic function is a parabola. What is the equation of the axis of symmetry? y=x^2-8x+16

OpenStudy (anonymous):

Put this in vertex form, i.e. \(\bf y=a(x-h)^2+k\) where \(\bf (x=h)\) is the equation of the axis of symmetry.

OpenStudy (anonymous):

watttttttttttttttttttttttttttttttttt

OpenStudy (anonymous):

do u know vertex form...

OpenStudy (anonymous):

to much work

OpenStudy (anonymous):

well there is another way..find the zeroes of the quadratic first. can you do that?

OpenStudy (anonymous):

here The graph of this quadratic function is a parabola. What is the equation of the axis of symmetry? y=x^2-3

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y%3Dx^2-3

OpenStudy (anonymous):

@beedhelpwithmath why did u change the question?

OpenStudy (anonymous):

cuz it was thw wrong one usin that link i sent u can i find the anwser

OpenStudy (anonymous):

@beedhelpwithmath just use wolframalpha to find the answer. wat are u going to do on a test? Think of it like this. When you graph a parabola, the axis of symmetry is the vertical line that divides the parabola into 2 halves.Here the axis of symmetry is x = 0 because it divides the parabola in to 2 SYMMETRICAL halves. Symmetry refers to exactly the same on both sides. Now, for a parabola, the vertex, which is either the peak of the parabola if it opens down or a minimum point of the parabola if it opens up, divides the parabola in to exactly 2 halves. It's x-coordinate is the axis of symmetry, i.e. if (a, b) is the vertex of a parabola then the axis of symmetry is x = a. This why it helps to convert from the standard y = ax^2 + bx + c form to y = a(x-h)^2 + k known as the vertex form because it tells you exactly what the vertex is, as a matter of fact the vertex will be at (h, k) and so the axis of symmetry will be x = h.